OpenStudy (anonymous):

Lisa ate 1/8 of a pizza. Josh ate 2 1/2 times a much as Lisa. What fraction of the pizza did Josh eat?

OpenStudy (aravindg):

2 1/2=5/2 fraction of Pizza Josh eat =(5/2) of ( 1/8) that will be?

OpenStudy (aravindg):

@bookaholic we at openstudy do not appreciate giving away answer like that ,instead help the asker get to the solution. Refer Openstudy COC http://openstudy.com/code-of-conduct.Future repitition of this can result in ban .Thank you

OpenStudy (anonymous):

Sorry @ AravindG